OpenStudy (anonymous):

I really need to get these questions right can someone help me check my work 1. which of the following is not an equivalent form of the compound inequality x-1< -11 or x+12>11 a. x<-10 or x> -1 b. a number line with an open circle on -10, shading to the left, and an open cirlce on -1, shading to the right c. x>-1 or x<-10 d. a number line with an open circle on -10, and open circle on -1, and shading in between. MY ANSWER: D?

OpenStudy (raden):

u are right

OpenStudy (anonymous):

thank you so much c: my grade is on the line so im trying to do the best I can but I'm not good at math so i appreciate it

OpenStudy (anonymous):

do you mind me asking one more question?

OpenStudy (raden):

what is it ?

OpenStudy (anonymous):

which of the following is not an equivalent form of the compound inequality x+16=<11 & x+16>= 3 a. 11 >= x +16>=3 b. -5>+=>=-13 c. a number line with a closed circle on -13, a closed circle on -5 and shading in between d. a number line with a closed circle on -13 shading to the left and a closed circle on -5 shading to the right i think its B

OpenStudy (raden):

i think D is wrong, so the answer is D

OpenStudy (anonymous):

okay thank you so much for helping me c:

OpenStudy (raden):

welcome